### v5

@@ 5.3 Marrying the upper and lower bounds
With this definition, we can rephrase Lemma 28 as stating simply that
${\mathbb{E}}[Z]\leq c_{3}q+1/2\cdot L^{2}$
We next derive a lower bound in terms of $q$ as well.
@@
From Lemma 31, we have ${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q=q\cdot\frac{\gamma{\mathbb{E}}[D]+c_{5}Q}{{\mathbb{E}}[D]/c_{3}+Q}\leq q\cdot\max\{\frac{\gamma}{1/c_{3}},\frac{c_{5}}{1}\}$; it can be checked numerically that $\max\{c_{5},\gamma c_{3}\}\leq c_{6}^{2}$. Now consider random variable $U=A^{2}$. The function $u\mapsto(L-\sqrt{u})^{2}$ is concave-up, so by Jensen’s inequality,
@@
Finally, the bound ${\mathbb{E}}[U]\leq c_{6}^{2}q$ implies that $(L-\sqrt{{\mathbb{E}}[U]})^{2}\geq\max\{0,L-c_{6}\sqrt{q}\}^{2}$. ∎
@@
By Corollary 19, we need to bound the ratio ${\mathbb{E}}[Z]/\operatorname{LB}$. Let $v=\sqrt{q}/L$. From Proposition 33, we have:
@@ Appendix A Proof of Lemma 14
Observation 36.
For real numbers $a_{1},a_{2},a_{1}^{\prime},a_{2}^{\prime},b$ with $a_{1}\geq a^{\prime}_{1}\geq 1,a_{2}\geq a^{\prime}_{2}\geq 1$ and $b\geq 0$, there holds $\frac{(a_{1}-1)(a_{2}-1)}{a_{1}a_{2}+b}\geq\frac{(a_{1}^{\prime}-1)(a_{2}^{\prime}-1)}{a^{\prime}_{1}a^{\prime}_{2}+b}$.
Proof.
It can shown mechanically e.g via decidability of first-order theory of real-closed fields. ∎
We are now ready to show Lemma 14. Rephrasing Theorem 13, we want to show that:
@@
Let $z=t^{2}/2-t$. By combining Proposition 35 and Observation 36, we get
@@ B.1 Calculation of $c_{1}$
$\frac{s^{2}\cdot\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}}{2(s-\kappa r)}=\frac{r(\mathrm{e}^{1/r}-1)}{2\mathrm{e}^{1/r}+2}\cdot\frac{t^{2}}{t-\kappa}\geq\frac{\theta(\mathrm{e}^{1/\theta}-1)}{2\mathrm{e}^{1/\theta}+2}\cdot 4\kappa\approx 0.591909465..$
@@
Since the RHS of Eq. (11) is an algebraic function of $r,s,x$, it can be checked automatically that its mimimum value is $0.6561$ over the domain $r\in[\theta,\tau],s\in[\max\{r,4/3\},\pi r],x\in[0,r]$. Putting the two cases together, we have shown that $c_{1}^{\prime}\geq 0.591909465..$.
@@
$c_{1}^{\prime\prime}:=\min_{\begin{subarray}{c}r\in[0,\theta],x\in[0,r]\\ y\in[\tau-r,1],s\in[r,\pi r]\\ d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl{(}x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+\frac{2d((1-x/\tau)^{1-1/x}-1)((r/\tau)^{1-1/y}-1)}{(1-x/\tau)^{1-1/x}(r/\tau)^{1-1/y}-(r-x)/\tau}\bigr{)}}{2(s+d-\kappa(r+y))}$ (12)
@@
We will divide the search region for $r,y$ into boxes $[r_{\min},r_{\max}]\times[y_{\min},y_{\max}]$ of width $\varepsilon=1/1000$. Within a given box (and subject to all other constraints on the variables), we can approximate:
@@
for $u_{0},u_{1},u_{2}$ rational numbers within $10^{-7}$ of $\mathrm{e}^{1/\tau},(r_{\max}/\tau)^{1-1/y_{\max}},\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$ respectively. By Observation 36 and these other approximations, we can lower-bound the expression in Eq. (12) by:
@@ B.2 Calculation of $c_{5}$
[…]
$\int_{h=1}^{\pi}\frac{F(h)\ dh}{h^{3}}\leq\sum_{i=0}^{\lceil(\pi-1)/\varepsilon\rceil-1}\int_{h=1+i\varepsilon}^{1+(i+1)\varepsilon}\frac{F^{\max}[1+i\varepsilon,1+(i+1)\varepsilon]\ dh}{h^{3}}$

### v8

@@ 5.3 Marrying the upper and lower bounds
Observation 31.
There holds ${\mathbb{E}}[Z]\leq c_{3}q+L^{2}/2$
Proof.
Restatement of Lemma 27. ∎
@@
From Lemma 30, we have ${\mathbb{E}}[A^{2}]\leq\gamma{\mathbb{E}}[D]+c_{5}Q=q\cdot\frac{\gamma{\mathbb{E}}[D]+c_{5}Q}{{\mathbb{E}}[D]/c_{3}+Q}\leq q\cdot\max\{\frac{\gamma}{1/c_{3}},\frac{c_{5}}{1}\}$; it can be checked numerically that this is at most $c_{6}^{2}q$. Now consider random variable $U=A^{2}$. The function $u\mapsto(L-\sqrt{u})^{2}$ is concave-up, so by Jensen’s inequality,
@@
The bound ${\mathbb{E}}[A^{2}]={\mathbb{E}}[U]\leq c_{6}^{2}q$ implies that $(L-\sqrt{{\mathbb{E}}[U]})^{2}\geq\max\{0,L-c_{6}\sqrt{q}\}^{2}$. ∎
@@
By Corollary 19, we need to bound the ratio ${\mathbb{E}}[Z]/\operatorname{LB}$. Let $v=\sqrt{q}/L$. From Observation 31 and Proposition 33, we have:
@@ Appendix A Proof of Lemma 14
Observation 35.
For $a_{1}\geq a^{\prime}_{1}\geq 1,a_{2}\geq a_{2}^{\prime}\geq 1,b\geq 0$, there holds
$\frac{(a_{1}-1)(a_{2}-1)}{a_{1}a_{2}+b}\geq\frac{(a_{1}^{\prime}-1)(a_{2}^{\prime}-1)}{a^{\prime}_{1}a^{\prime}_{2}+b}.$
Proof.
Can be shown mechanically e.g via decidability of first-order theory of real-closed fields. ∎
@@
We now turn to proving Lemma 14. Rephrasing Theorem 13, we want to show that:
@@
Let $z=t^{2}/2-t$. By combining Observation 35 and Proposition 36, we get
@@ B.1 Calculation of $c_{1}$
$\frac{s^{2}\cdot\frac{\mathrm{e}^{1/r}-1}{\mathrm{e}^{1/r}+1}}{2(s-\kappa r)}=\frac{r(\mathrm{e}^{1/r}-1)}{2\mathrm{e}^{1/r}+2}\cdot\frac{t^{2}}{t-\kappa}\geq\frac{\theta(\mathrm{e}^{1/\theta}-1)}{2\mathrm{e}^{1/\theta}+2}\cdot 4\kappa\approx 0.591909465$...,
@@
Since the RHS of Eq. (12) is an algebraic function of $r,s,x$, it can be checked automatically that its mimimum value is $0.6561$... over the domain $r\in[\theta,\tau],s\in[\max\{r,4/3\},\pi r],x\in[0,r]$. Putting the two cases together, we have shown that $c_{1}^{\prime}\geq 0.591909465..$.
@@
$c_{1}^{\prime\prime}:=\inf_{\begin{subarray}{c}r\in[0,\theta],x\in(0,r]\\ y\in[\tau-r,1],s\in[r,\pi r]\\ d\in[y,\pi y]\end{subarray}}\frac{d^{2}+s\bigl(x+(\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}-c_{0}x)(s-x)+2d\Phi(x,y;x/\tau,1-r/\tau)}{2(s+d-\kappa(r+y))}$ (13)
where
$\Phi(x,y;x/\tau,1-r/\tau)=\frac{((1-x/\tau)^{1-1/x}-1)((r/\tau)^{1-1/y}-1)}{(1-x/\tau)^{1-1/x}(r/\tau)^{1-1/y}-(r-x)/\tau}.$
@@
We will divide the search region for $r,y$ into boxes $r_{\min},r_{\max}]\times[y_{\min},y_{\max}]$ of width $\varepsilon=1/1000$. Within a given box (and subject to all other constraints on the variables), we can approximate:
@@
for $u_{0},u_{1},u_{2}$ rational numbers within $10^{-7}$ of $\mathrm{e}^{1/\tau},(r_{\max}/\tau)^{1-1/y_{\max}},\frac{\mathrm{e}^{1/\tau}-1}{\mathrm{e}^{1/\tau}+1}$ respectively. By Observation 35 and these other approximations, we can lower-bound the expression in Eq. (13) by:
@@ B.2 Calculation of $c_{5}$
This can be evaluated easily by Mathematica for given $h_{\min},h_{\max}$. Note that the constraint $r\in[0,\theta]$ has been relaxed to $r\in\mathbb{R}$; this is necessary for Mathematica to compute the optimization efficiently. To handle the overall integral $\int_{h=1}^{\pi}F(h)/h^{3}\ \mathrm{d}h$, we divide the integration region $1,\pi]$ into strips of width $\varepsilon=10^{-3}$, and estimate:
$\int_{h=1}^{\pi}\frac{F(h)\ \mathrm{d}h}{h^{3}}\leq\sum_{i=0}^{\lceil(\pi-1)/\varepsilon\rceil-1}\int_{h=1+i\varepsilon}^{1+(i+1)\varepsilon}\frac{F^{\max}[1+i\varepsilon,1+(i+1)\varepsilon]\ \mathrm{d}h}{h^{3}}$